1. A subject’s measured volumes and capacities are as follows:

a. TLC = 5500 ml

b. IRV = 2200 ml

c. ERV = 1500 ml

d. VT = 650 ml

i. What is his FRC in ml?

1. 4000

2. 2650

3. 2300

4. Can’t tell from the available data
Answer #1

ANSWER: The correct option is 2) 2650 ml and the calculations are as follows:

FRC is Functional residual capacity and is calculated by the following formula:

FRC=RV + ERV, where RV is residual volume & ERV is expiratory reserve volume.

In the given question, we have got the following values:

  • TLC (Total Lung Capacity) =5500mL
  • IRV (Inspiratory Reserve volume)= 2200mL
  • ERV= 1500mL
  • VT (Tidal Volume)= 650mL

The formula to calculate TLC= TV+ IRV+ ERV+ RV

Hence, RV= TLC-(TV+IRV+ERV)
=5500- (650+2200+1500)
=5500- 4350
=1150 mL
Here we got RV as 1150 ml.

Now, we can calculate FRC=RV + ERV
=1150 + 1500
= 2650 ml

Hence, we got FRC as 2650 ml. So, the correct option is 2.